write an equation of the line containing the given point and paralled to the given line (-4,3) ;4x=9y+8
Thank you so very much. You just don't know how much I appreciate your help. Thank you, again!

Answer by scott8148(6628)

parallel means that the slope is the same , so the x and y coefficients don't change

the new line will have a different constant term (the +8)
___ you can find the new constant by plugging in the given point

4(-4) = 9(3) + k ___ -16 = 27 + k ___ -43 = k

the equation for the new line is ___ 4x = 9y - 43